6x+3y=9 x=0 y=? x=2 y=?
Answer by nyc_function(2741)


When x = 0, we solve for y.
6x + 3y = 9
6(0) + 3y = 9
0 + 3y = 9
3y = 9
y = 9/3
y = 3
The first point to graph is (0, 3).
====================================
When x = 2, solve for y.
6x + 3y = 9
6(2) + 3y = 9
12 + 3y = 9
3y = 9 - 12
3y = -3
y = -3/3
y = -1
The second point to graph is (2, -1).
Graph each point on the coordinate plane and connect the points with a straight line.